A fence of length 50 meters has posts that are 2 meters apart. How many posts would there be in the 50 meter fence?
Can you please help me with this question?

Answer by stanbon(75887)

A fence of length 50 meters has posts that are 2 meters apart. How many posts would there be in the 50 meter fence?
----------------------------------------------------
A fence of 2 meters would need 2 posts; 2/2 = 1
A fence of 4 meters would need 3 posts: 4/2 = 2
.....
A fence of 10 meters would need 6 posts: 10/2 = 5
So,
A fence of 50 meters would need 26 posts: 50/2 = 25
==================
Cheers,
San H.